OpenStudy (anonymous):

what is 5 2/3 divided by 2 1/6

OpenStudy (diyadiya):

17/3 / 13/6=17/3*6/13=34/13

OpenStudy (anonymous):

thanks

OpenStudy (diyadiya):

which is equal to 2 8/13

OpenStudy (anonymous):

\[2\frac{8}{13} \]